OpenStudy (deercult):

A candle is 17 in. tall after burning for 3 hours. After 5 hours, it is 15 in. tall. Write a linear equation to model the relationship between height 'h' of the candle and time 't'. Predict how tall the candle will be after burning 8 hours. --- I'd like to attempt the last part on my own. I'm just having trouble writing a linear equation. So far: -I understand that it is burning at one inch an hour. For example, it was probably 16 inches at 4 hours. I gathered this from simple math (it has been two hours, and it has melted two inches.) -I understand that there's a negative involved.

OpenStudy (deercult):

I don't have any ideas. I've had two but my gut tells me they're wrong. I'll provide if you want, though, but again, they're almost certainly incorrect.

OpenStudy (deercult):

@mathstudent55 (If you aren't busy, of course)

OpenStudy (wckeller):

Oh, I was going to suggest using the slope formula and point-slope formula (using the hours and height as x and y) to find it.

OpenStudy (mathstudent55):

Use the given info as two points, and write the equation of the line between them.

OpenStudy (mathstudent55):

3 hours, 17 in. 5 hours, 15 in. Use the points (3, 17) (5, 15)

OpenStudy (deercult):

Alright, so thatd be \[m = \frac{ 17-15 }{ 3-5 }\] Right?

OpenStudy (deercult):

Just making sure, I don't memorize formulas well.

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Good.

OpenStudy (mathstudent55):

Did you find the slope?

OpenStudy (deercult):

Which leaves me with a slope of \[\frac{ 2 }{ -2 }\]

OpenStudy (mathstudent55):

Correct, which reduces to?

OpenStudy (deercult):

-1, yes?

OpenStudy (mathstudent55):

Correct. Now you need to use the slope and one point to find the equation of the line.

OpenStudy (deercult):

\[y - 17 = -1(0 - 3)\] Is this valid for finding the y-intercept for the equation??

OpenStudy (mathstudent55):

I'd do this: y = mx + b Now plug in one of the known points into x and y, and plug in the slope into m. Let's use point (3, 17) 17 = -1*3 + b Now we can find the y-intercept, b.

OpenStudy (deercult):

Ohh. That way I got -20. The way I did it (using the slope-point formula) I got 20. Judging by the graph (+ it'd make no sense in the context of a candle burning and all, esp. with the numbers given) I think it's the positive one.

OpenStudy (mathstudent55):

I used the slope-intercept equation above. You can also use the point-slope equation: \(y - y_1 = m(x - x_1)\) \(y - 17 = -1(x - 3) \) You will see that they will give you the same equation.

OpenStudy (mathstudent55):

b = 20

OpenStudy (deercult):

Leaving y = -1x + 20

OpenStudy (mathstudent55):

Correct y = -x + 20 The want h for the height and t for time, so let's change to their variables: h = -t + 20

OpenStudy (mathstudent55):

As a quick check, let's make sure the given points work. Check (3, 17) h = -t + 20 17 = -3 + 20 17 = 17 Check (5, 15) h = -t + 20 15 = -5 + 20 15 = 15 Yes, both points do work.

OpenStudy (mathstudent55):

Now use the equation we got, and do the last part. Here, you need time = 8, so let t = 8 in our equation h = -t + 20

OpenStudy (deercult):

-8 + 20 = 12 Meaning after 8 hours of burning it's 12 inches tall.

OpenStudy (mathstudent55):